OpenStudy (anonymous):

F=(9/5)C+32 will someone please help me

OpenStudy (anonymous):

is this algebra 1??

OpenStudy (anonymous):

okay just hold on i gotcha (:

OpenStudy (anonymous):

when fahrenheit temp is 50, just plug in 50 in the place of F in your formula. Therefore, 50 = (9/5) C + 32 (9/5) C = 50 - 32 = 18 C = 18*5/9 = 10 Therefore, temperature in 10 degree celcius does that help you its prety much the whole answer (: and youre welcome